OpenStudy (slipknot306):

Combine the like terms to simplify the expression: 9 + 3 + 17xy2 + 8y3 + 10x –13xy2 – 9x – 6y3 A. 2y6 + 4x2y2 + 10x + 12 B. 2y3 – 4xy2 – 3x + 12 C. 2y3 + 4xy2 + x + 12 D. 2y3 + 4xy2 + 10x + 9

OpenStudy (texaschic101):

Like terms have exactly the same letters... 9 + 3 + 17xy^2 + 8y^3 + 10x - 13xy^2 - 9x - 6y^3 so lets combine like terms.. 17xy^2 - 13xy^2 = 4xy^2 8y^3 - 6y^3 = 2y^3 10x - 9x = x 9 + 3 = 12 so we have : 4xy^2 + 2y^3 + x + 12...re-arranged and it would be C